OpenStudy (anonymous):

If the sin 45° is 0.707, then the cos _____° is 0.707.

OpenStudy (anonymous):

Is it 45 degrees also? Idk

OpenStudy (anonymous):

@mathmale @mertsj

OpenStudy (raden):

use these identities : sin(x) = cos(90 - x) also sin(x) = cos(270 + x) here plug in that x = 45